### v1

# Second-order Optimization of Gaussian Splats with Importance Sampling

Abstract
[…]
Project Page: https://vcai.mpi-inf.mpg.de/projects/LM-IS
1 Introduction
Figure 1: We introduce the first second-order optimizer for Gaussian splatting. Notably, our dedicated optimizer converges significantly faster than Adam [23] and already achieves reasonable renderings after very few seconds of training.
Photoreal novel view synthesis from multi-view images or video has attracted significant attention in recent years due to widely applicable downstream tasks in content creation, VR/XR, gaming, and the movie industry, to only name a few. Here, Neural Radiance Fields (NeRF) [35] and 3D Gaussian Splatting (3DGS) [21] mark a major milestone, due to their unprecedented quality leading to follow ups beyond view synthesis like VR rendering [53], avatar creation [49], simultaneous localization and mapping (SLAM) [34, 1], and scene editing [3, 10].
However, NeRF-based models often require substantial training time, and researchers have developed various techniques to mitigate this problem. Some of them include neural hashing [38], employing explicit scene modeling [46], improved sampling strategies [19], and tensor factorization methods [9]. 3DGS [21] instead does not rely on coordinate-based representations, but leverages a set of 3D Gaussians, which can be effectively rendered into image space using tile-based rasterization. Nonetheless, optimizing the parameters of each Gaussian can still take hours. Previous work focused on finding better densification strategies [30, 22], quantization and compression of Gaussians [15, 40], and more efficient implementations for the backward pass [30], which led to substantially reduced training times. However, all of them mostly rely on a first-order optimization routine, i.e. gradient descent or Adam [23], and do not explore other (potentially second-order) alternatives.
Second-order optimization is known for having better convergence guarantees compared to first-order methods. However, its adoption in 3DGS remains challenging due to the high memory and computational demands associated with storing and inverting large Jacobian matrices. Notably, the Jacobian size scales linearly with both the number of Gaussians and image pixels. Unlike NeRFs, which rely on a dense neural network and a coordinate-based formulation, 3DGS benefits from inherently sparse Jacobians since each Gaussian influences only a small subset of pixels residuals. Our key idea is to leverage this sparsity in order to make second-order optimization for Gaussian splatting not only tractable, but also potentially more efficient than first-order baselines.
To this end, we formulate fitting the Gaussian parameters to the multi-view images as a non-linear least squares optimization problem and leverage the Levenberg-Marquardt (LM) algorithm for solving it. Yet, adopting the LM algorithm poses significant challenges due to the high storage and computation requirements of the Jacobian matrix. Furthermore, solving the numerical system arising from the normal equation is computationally expensive, potentially limiting feasibility at larger scales.
To overcome these challenges, we propose a GPU-parallelized conjugate gradient solver for our second-order 3DGS optimizer, which circumvents explicit storage of the Jacobian matrix, and solves the matrix inverse iteratively. Firstly, we show that implementing the solver naively does not result in a fast optimizer because of the high computational cost of Jacobian-vector products needed in the conjugate gradient solver. Therefore, we propose to approximate the full normal equation by an effective view sampling strategy and by importance sampling individual pixels, which results in significantly faster convergence. We also introduce a heuristic to automatically determine the learning rate, which eliminates the need for line search algorithms that are commonly used in conjunction with second-order optimizers.
Overall, with these improvements, our proposed second-order optimizer is both memory and compute efficient, suggesting that second-order optimization for 3DGS is a highly promising direction for further study. Our optimizer demonstrates significant improvements in settings with a low number of Gaussians, and can compete with first-order optimizers when the number of Gaussians is moderate, all without explicitly storing the Jacobian. In summary, we propose a novel second-order optimizer for 3DGS with the following features:
  • Formulating Gaussian splatting as a non-linear least squares optimization problem that is solved using a memory and computationally efficient Levenberg-Marquardt and conjugate gradient solver specifically tailored towards 3DGS.
  • A view and importance sampling strategy over the pixels (residuals) to effectively approximate the loss, leading to a significant decrease in computational complexity.
  • An effective heuristic to determine the learning rate, which eliminates the need for expensive line search methods while providing stable convergence for 3DGS optimization.
2 Related Work
3D Scene Representation. To model 3D scene representations, Neural Radiance Fields (NeRFs) [35] represent the 3D geometry of a scene implicitly using a coordinate-based neural network. Subsequently, several variations of NeRFs have been proposed to improve visual quality [4] and to improve computational efficiency [48, 17, 39]. However, despite the significant advancements, NeRFs still face limitations, particularly in terms of their training and rendering efficiency. More recently, 3D Gaussian Splatting [21] has been proposed, which employs a fully explicit representation, unlocking significant enhancements in rendering speed while maintaining similar qualitative results to NeRF-based methods. Recent advancements have further improved the rendering quality of 3DGS [29, 55], or its efficiency [13, 27, 14, 43, 30]. Notably, LightGaussian [13] and C3DGS [27] focus on pruning Gaussians, directly leading to improvements in rendering speed. Another related approach is to constrain the densification process, efficiently modeling the scene with less Gaussians, such as in Mini-Splatting [14] and 3DGS-MCMC [22]. Some other works aim to compress or quantize Gaussians [15, 40]. Taming3DGS [30] steers the densification process in a controlled manner, while also devising a parallelized implementation for more efficient backpropagation in CUDA. Orthogonally to these works, we explore second-order optimizers for 3DGS optimization, which is challenging due to the high memory and computational demands associated with storing and inverting large Jacobian matrices. To achieve this, we exploit the sparsity of the Jacobian matrices of 3DGS.
First-order Optimizers. Stochastic Gradient Descent (SGD) [42] was developed as a general optimization method and has been widely used in modern multi-dimensional deep learning tasks. With the addition of momentum [45], SGD can escape from local optima, a crucial feature in stochastic settings. However, a key limitation is that SGD applies the same update to all dimensions, which makes it difficult to adapt to complex problems. To address this, researchers have developed various preconditioners for the gradient information. Adagrad [12] accumulates the squares of the gradients and preconditions the gradient with the square root of the accumulated values. RMSprop [50] replaces accumulation with an exponentially moving average, which leads to more stable training. Adam [23], arguably the most popular optimizer today, combines momentum with the moving average of second-order moments. It has also become the default optimizer for many 3D tasks, including NeRF and 3DGS, and provides a strong baseline. However, despite the efficacy of first-order optimizers, there is strong incentive to explore second-order optimizers due to their attractive properties, as discussed next.
Second-order Optimizers. Second-order optimizers often approximate the inverse of the Hessian matrix to precondition the solution, and can offer significant advantages over first-order methods. They require far fewer iterations because they can obtain the second-order approximation of the loss landscape into account [32, 31, 33]. Additionally, second-order methods are often not as sensitive to learning rate values, and can often estimate it locally through line search algorithms [2, 18, 37], or trust region methods [11, 41].
A subset of second-order optimizers, particularly those deriving from Gauss-Newton, leverage curvature information to refine parameter updates. The general update rule for this family of optimizers follows:
$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\mathbf{H^{-1}}\mathbf{g}$ (1)
where $\mathbf{g}\in\mathbb{R}^{n}$ is the gradient, $\mathbf{H}\in\mathbb{R}^{n\times n}$ is the Hessian of the loss function and $\boldsymbol{\beta}\in\mathbb{R}^{n}$ is the update step. In a naive implementation with $n$ parameters, storing the Hessian requires $O(n^{2})$ space, and computing its inverse takes $O(n^{3})$ time, making it impractical for large-scale optimization tasks. Instead, a numerical approximation of the inverse can be obtained using iterative solvers like conjugate gradient (CG). This method also eliminates the need to store the Hessian matrix, as CG only requires the results of matrix-vector multiplications [31, 33]. In the literature, this class of methods is referred to as Hessian-free or matrix-free optimization.
Specifically, within the class of second-order optimizers, the Gauss-Newton (GN) algorithm [26] is sometimes adopted to approximate the hessian $\mathbf{H}$ with $\mathbf{J^{\top}J}$, where $\mathbf{J}$ is the Jacobian matrix. This ensures that the Hessian’s approximations are positive semi-definite, which guarantees the existence of a solution to the normal equation and avoids local minima. Notably, the Levenberg-Marquardt (LM) algorithm [36] is an extension over GN, interpolating between GN and gradient descent, resulting in more stable optimization. Notably, some concurrent works [25, 18] also explore second-order optimizers for 3DGS optimization. 3DGS 2 [25] explores an algorithm based on Newton’s method, but makes it computationally tractable by limiting the second-order computations to a small locality. Concurrently, 3DGS-LM [18] adopts the LM optimizer, where they propose a caching data structure to store intermediate gradients for fast computation.
In our paper, we adapt the LM optimizer, proposing a fast GPU-parallelized conjugate solver for 3DGS, which is a matrix-free optimization that avoids explicitly storing the Jacobian matrix, and solves the matrix inverse iteratively. Our LM optimizer exploits the sparsity property within 3DGS representations for further efficiency by an effective view sampling strategy and by importance sampling individual pixels (residuals), which results in significantly faster convergence. We also create a heuristic to automatically determine the learning rate, which eliminates the need for line search algorithms.
Figure 2: Overview of our method is given. We start from randomly initialized Gaussians and gradually refine them with Levenberg-Marquardt optimizer. Since dealing with the true Jacobian matrix is costly, we approximate it with a tile-aware sampling algorithm. After we solve the normal equations with approximated Jacobians, we update the parameters using a learning rate heuristic. Note that the Jacobians are never materialized in the memory, and the normal equation is solved with only Jacobian vector products.
3 Background - 3D Gaussian Splatting
3DGS is a point-based representation that can reconstruct 3D scenes with high fidelity. The 3D scene is represented using a set of Gaussians, where each Gaussian has an optimizable mean vector $\mathbf{x}$ and covariance matrix $\Sigma$:
$G(\mathbf{x})=e^{-\frac{1}{2}\mathbf{x}^{T}\boldsymbol{\Sigma}^{-1}\mathbf{x}}$ (2)
To ensure the covariance matrix remains positive-definite, it is decomposed into rotation $\mathbf{R}$ and scale $\mathbf{S}$ matrices:
$\Sigma=\mathbf{R}\mathbf{S}\mathbf{S^{\top}}\mathbf{R^{\top}}.$ (3)
To render a pixel, the Gaussians are first projected onto the 2D image plane, and sorted according to their depth. A pixel value $C_{i}$ is computed using $\alpha$-blending, which combines the color $c$ and per-pixel opacity $\alpha$ of the projected Gaussians:
$C_{i}=$ $\sum_{n\leq\mathcal{S}}c_{n}\cdot\alpha_{n}\cdot\prod_{m<n}(1-\alpha_{m})$ (4)
[…]
where $\mathcal{S}$ represents the total number of depth-sorted Gaussians projected onto pixel $i$, and each $\alpha_{n}$ is obtained by multiplying its corresponding Gaussian opacity $o_{n}$ and 2D projection (via camera projection $\boldsymbol{\pi}_{\text{cam}}$).
The parameters of the Gaussians are updated by optimizing a loss function between the rendered image $\hat{\mathcal{I}}$ and ground-truth image ${\mathcal{I}}$. This optimization is done predominantly using the Adam optimizer:
$\mathcal{L}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{f}(\hat{\mathcal{I}},\mathcal{I}),$ (5)
where $\mathcal{L}_{f}(\cdot)$ is a loss function that quantifies image differences. In this paper, we experiment with mean squared error (MSE) and structural similarity (SSIM) [51] loss functions. 3DGS models the view-dependent effects with spherical harmonics. In this work, however, we assume each scene has Lambertian surface and disable the corresponding spherical harmonics levels.
4 Method
We first introduce the LM optimizer in Sec. 4.1, which we adopt in this work. Then, we derive why a naive implementation is not feasible for Gaussian splatting. To resolve this issue, in Sec. 4.2, we discuss our matrix-free approach to adapt the LM optimizer. In detail, we propose a GPU-parallelized conjugate gradient solver for our second-order 3DGS optimizer, which circumvents explicit storage of the Jacobian matrix, and solves the matrix inverse iteratively. Next, in Sec. 4.3, we introduce a new view sampling strategy to effectively approximate the full normal equation, thereby providing reliable update step directions by integrating information from multiple views. In Sec. 4.4, we present our importance sampling of individual pixels (residuals), providing an approximate loss function, which results in significantly faster convergence. Lastly, in Sec. 4.5, we introduce a heuristic to automatically determine the learning rate, which eliminates the need for line search algorithms that are commonly used in conjunction with second-order optimizers.
4.1 Levenberg-Marquardt Optimizer for 3DGS
We use the LM optimizer to compute an update step by minimizing the loss over a mini-batch $B$, which includes images of the shape height $H$, width $W$, and channels $C$. The optimizable parameters for each Gaussian are opacity $o\in\mathbb{R}$, color $c\in\mathbb{R}^{3}$, mean value $\mathcal{X}\in\mathbb{R}^{3}$ , scale $s\in\mathbb{R}^{3}$ , and quaternion rotation $q\in\mathbb{R}^{4}$. We represent all Gaussian parameters with $\boldsymbol{\beta}\in\mathbb{R}^{P}$, where we denote the total number of optimizable parameters with $P$.
The rendering loss function in Eq. 5 can be rewritten as a nonlinear least squares objective:
$\mathcal{L}=\sum_{i=1}^{M}(r_{i})^{2}$ (6)
where $r$ is a residual and $M=BHWC$. The residuals are defined as pixel and structural similarity losses:
$r_{i}=\mathcal{I}_{i}-\hat{\mathcal{I}_{i}}$ (7)
Then, the update vector $\Delta\boldsymbol{\beta}\in\mathbb{R}^{P}$ is retrieved by solving the following normal equation:
$\left(\mathbf{J}^{T}\mathbf{J}+\lambda\mathbf{I}\right)\Delta\boldsymbol{\beta}=-\mathbf{J}^{T}\mathbf{r}$ (8)
where $\lambda$ is the damping parameter, and $\mathbf{J}\in\mathbb{R}^{M\times P}$ is the Jacobian of $\mathbf{r}$.
One iteration of the LM optimizer is completed after we update all the Gaussian parameters $\boldsymbol{\beta}$ with the learning rate $\eta$:
$\boldsymbol{\beta}_{(k+1)}=\boldsymbol{\beta}_{(k)}+\eta\cdot\Delta\boldsymbol{\beta}$ (9)
Note that in the original 3DGS work, for each parameter group (opacity, color, mean, scale, and rotation), a different learning rate is used. On the other hand, we use a uniform learning rate across parameters because Gauss-Newton-type methods inherently incorporate parameter scaling through the Hessian approximation $\mathbf{J^{\top}J}$. In Sec. 4.5, we discuss a heuristic to determine the learning rate at every iteration.
In practice, naively solving the normal equation is not feasible, especially in large-scale numerical systems. First, the Jacobians scale linearly with both the number of Gaussians and the number of pixels, and storing the Jacobians explicitly in memory quickly becomes infeasible. For example, using $100$ images at a resolution of $800\times 800$ pixels as training data, along with 10 0001000010\,00010 000 Gaussians to model the 3D scene, would result in Jacobians exceeding $100$ TB in size. Even assuming a sparse storage format with a $99\%$ sparsity, the storage requirement would still exceed $1$ TB for a true LM optimizer. In addition, taking the inverse of the left-hand side is computationally expensive, with O($P^{3}$) time complexity when implemented naively. Thus, computing the explicit inverse is often infeasible, and previous work has focused on solving the normal equation iteratively [31, 33, 8], but they still remain computationally expensive, as discussed in the next section.
4.2 Matrix-free Preconditioned Conjugate Gradient (PCG) Solver for 3DGS
We first discuss our matrix-free PCG solver to adapt the LM optimizer. In this subsection, we present a GPU-parallelized conjugate gradient solver, which does not need to store the Jacobian matrix explicitly, avoiding the aforementioned memory issues. To achieve this, we solve Eq. 8 via the preconditioned conjugate gradient algorithm, which only needs results of Jacobian-vector products, and finds the solution $\Delta\boldsymbol{\beta}$ iteratively. Note that while explicit storage of certain Jacobian elements enhances the performance of the PCG solver as demonstrated in [18], our work prioritizes a fully matrix-free implementation that minimizes memory usage.
We choose the Jacobi preconditioner in our implementation due to its simplicity and effectiveness. More specifically, we use $\frac{1}{\operatorname{diag}(\mathbf{J^{\top}J}+\lambda\mathbf{I})}\in\mathbb{R}^{P}$, as the precondition. We provide the implementation of the solver in CUDA, together with efficient Jacobian vector product kernels. See the supplementary material for details about the kernel design. The pseudocode of the optimizer is given in Alg. 1.
Algorithm 1 One step of LM optimizer with matrix-free PCG solver.
Input: Gaussians $\boldsymbol{\beta}_{k}$, cameras $\mathcal{C}$
Output: Updated Gaussians $\boldsymbol{\beta}_{k+1}$
1:$\mathcal{I},\mathcal{C}_{B}=\text{getBatch}(\mathcal{C})$triangleright Sec. 4.3
2:^$\hat{\mathcal{I}}=\text{splatting}({\boldsymbol{\beta}_{k}},\mathcal{C}_{B})$
3:$\mathcal{\mathbf{r}}=\text{getResiduals}(\mathcal{I},\hat{\mathcal{I}})$triangleright Eq. 7
4$\triangleright$ Beginning of matrix-free PCG solver (Sec 4.2)
5:$\mathbf{r}_{0}=\mathbf{J^{\top}r}$triangleright Estimated in Sec. 4.4
6:$\mathbf{M}^{-1}=1/\text{Diag}(\mathbf{J^{\top}J}+\lambda\mathbf{I})$triangleright Estimated in Sec. 4.4
7:$\mathbf{z}_{0}=\mathbf{M}^{-1}\mathbf{r}_{0}$
8:$\mathbf{p}_{0}=\mathbf{z}_{0}$
9:$\mathbf{x}_{0}=\mathbf{0}$
10:for $i=0$ to $\text{PCG}_{\text{iters}}$ do
11:    $\mathbf{u}=(\mathbf{J^{\top}J}+\lambda\mathbf{I})\mathbf{p}_{i}$triangleright Estimated in Sec. 4.4
12:    $\alpha=\frac{\mathbf{r}_{i}^{T}\mathbf{z}_{i}}{\mathbf{p}_{i}^{T}\mathbf{u}}$
13:    $\mathbf{x}_{i+1}=\mathbf{x}_{i}+\alpha\mathbf{p}_{i}$
14:    $\mathbf{r}_{i+1}=\mathbf{r}_{i}-\alpha\mathbf{u}$
15:    $\mathbf{z}_{i+1}=\mathbf{M}^{-1}\mathbf{r}_{i+1}$
16:    $\beta=\frac{\mathbf{r}_{i+1}^{T}\mathbf{z}_{i+1}}{\mathbf{r}_{i}^{T}\mathbf{z}_{i}}$
17:    $\mathbf{p}_{i+1}=\mathbf{z}_{i+1}+\beta\mathbf{p}_{i}$
18:end for
19:$\boldsymbol{\beta}_{k+1}=\boldsymbol{\beta}_{k}+\eta\,x_{i+1}$triangleright LR Heuristic Sec. 4.5
20:return $\textbf{return }\boldsymbol{\beta}_{k+1}$
Although this optimizer is able to converge to the final solution in a very limited number of steps, it is 3×3$\times$3 × slower than our final method, as shown in Table 3. The main reason for this is that the conjugate gradient algorithm needs to run several iterations, therefore, we need to compute the $\mathbf{J}^{\top}\mathbf{J}\mathbf{v}$ product from Eq. 8 multiple times. This kind of computational complexity commonly arises when a second-order optimizer is used in large-scale numerical systems. The common approximations methods are diagonal [5, 28, 54, 47] and block-diagonal [32, 44, 6] approximation of the Hessian or $\mathbf{J^{\top}J}$ matrix.
We observe that the Jacobians in 3DGS do not exhibit a dominant diagonal or block-diagonal structure, as illustrated in Fig. 3. This arises because each pixel is rendered through the interaction of multiple Gaussians, as described in Eq. 4. In other words, no single Gaussian independently represents a surface; instead, it relies on the collective contribution of surrounding Gaussians to accurately capture the ground truth. This leads to many off-diagonal and off-block-diagonal entries in $\mathbf{J^{\top}J}$. Therefore, the common diagonal and block-diagonal approximations do not provide good approximations. To this end, we propose to estimate the loss function with our proposed view and importance sampling.
Figure 3: We visualize a normalized $\mathbf{J^{\top}J}$ matrix for one downsampled training image (a). While elements in the diagonal are common, we compute the dominance ratio as the normalized ratio between the diagonal element over the sum of other elements in the same row (b) and show how only a limited number of parameters lead to a diagonally dominant linear system (c).
4.3 View Sampling
Second-order methods like LM are typically used in deterministic settings where the full objective is evaluated at every iteration. If this is not the case, estimating local curvature can be problematic and become unreliable [8]. This poses a challenge in the 3DGS setting, where the number of views can exceed the hundreds, as incorporating all of them at the same time is infeasible. Yet, to compute meaningful gradients, we must find an effective way to approximate the full normal equation in Eq. 8.
To this end, we introduce a view sampling approach that allows us to get a diverse set of views in each batch. We run K-Means clustering to partition the camera locations into batch size number of clusters (e.g., 8 clusters created for a batch size of 8). An image is then randomly selected from every cluster. This approach ensures that each batch captures a more balanced and diverse set of views, resulting in a more accurate estimation of the curvature information. As evidenced by Table 2, this method converges to higher scores compared to the random sampling of the cameras.
4.4 Estimation of Loss Function with Importance Sampling
Additionally, to improve the efficiency and feasibility of our second-order optimizer for 3DGS, we further exploit the sparsity property within the 3DGS representation, which has been discussed in Sec. 4.2. Yet, at the same time, as observed in Fig. 3, there does not exist a well-organized structure to the sparsity. Instead, in this paper we propose to adopt an importance sampling approach, where we focus more on the important elements that give meaningful gradients.
We begin our derivation by rewriting our least squares loss function in Eq. 5 as an integral. Then, the loss function is given as:
[…]
where the integral is defined over $\Omega$, the union of all image domains of all cameras, and calculated at the coordinates $\mathbf{p}\in\mathbb{R}^{2}$ of the image plane. We also define $\ell(\mathbf{p})=r(\mathbf{p})^{2}$ to remove the squared term, which simplifies the subsequent derivations. To estimate the loss function $L$, we express it as an expectation with respect to an arbitrary probability distribution $q(\mathbf{p})$:
[…]
This formulation allows us to interpret $q(\mathbf{p})$ as a sampling distribution over the domain $\Omega$. Consequently, we can approximate the expectation numerically using Monte Carlo estimation. In particular, by independently sampling $N$ pixels $\{\mathbf{p}_{i}\}_{i=1}^{N}$ from $q(\mathbf{p})$, the loss function can be approximated as:
$L\approx\hat{L}=\frac{1}{N}\sum_{i=1}^{N}\ell(i)\,\frac{1}{q(i)}.$ (12)
The estimated loss function allows us to derive the estimated versions of Jacobian-vector products arising in the PCG solver. The derivative of the estimated loss function with respect to a parameter becomes:
$\frac{\partial\hat{L}}{\partial\boldsymbol{\beta}_{j}}=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial\ell(i)}{\partial\boldsymbol{\beta}_{j}}\,\frac{1}{q(i)}=\frac{2}{N}\sum_{i=1}^{N}\frac{\partial r_{i}}{\partial\beta_{j}}r_{i}\,\frac{1}{q(i)}$ (13)
This gradient corresponds to the right-hand side of the normal equation given in Eq. 8 since $\frac{\partial r_{i}}{\partial\beta_{j}}=J_{ij}$.
Similarly, we can calculate the Hessian of the estimated loss function given in Eq. 12. To do so, we take the derivative of Eq. 13 with respect to parameter $\boldsymbol{\beta}_{k}$:
$\hat{H}_{jk}=\frac{2}{N}\sum_{i=1}^{N}\left(\frac{\partial r_{i}}{\partial\beta_{j}}\frac{\partial r_{i}}{\partial\beta_{k}}\;+\;r_{i}\,\frac{\partial^{2}r_{i}}{\partial\beta_{j}\,\partial\beta_{k}}\right)\,\frac{1}{q(i)}$ (14)
When we apply the Gauss-Newton approximation to Hessian matrix, the term with second derivative in Eq. 14 is ignored, and we are left with an estimation of $\mathbf{J^{\top}J}$.
Note that this approximation preserves the symmetry and positive semi-definiteness of $\mathbf{J^{\top}J}$, a property required for the convergence of the conjugate gradient algorithm.
In this work, we experimented with simple uniform distribution $q_{\text{uniform}}(\mathbf{r})$ and a loss-based distribution $q_{\text{loss}}(\mathbf{r})$ to sample the pixels. Our intuition is that prioritizing pixels with high loss values facilitates the correction of the Gaussians responsible for these discrepancies. We define the likelihood of pixel $i$ using Softmax normalization as:
$q_{\text{loss}}(i)=Softmax(\|r_{i}\|)=\frac{\exp(\|r_{i}\|)}{\sum_{j}^{N}\exp(\|r_{j}\|)}$ (15)
Comparisons between the uniform and loss-based distribution are given in Table 3.
In practice, sampling pixels over the image does not efficiently integrate with GPU programming and the backward pass of 3DGS. The reason is that an image is divided into $16\times 16$ tiles and when the pixels are selected randomly, some tiles get more samples than others, causing an unbalanced workload among thread blocks. Moreover, the random nature of the sampling does not allow fixed thread assignment per tile. Therefore, we use a stratified sampling strategy by distributing samples among tiles and perform importance sampling inside them. In other words, instead of sampling $N$ pixels from the entire image, we sample $N/\mathcal{T}$ pixels from each tile, where $\mathcal{T}$ is the total number of tiles in the image.
The sampling mechanism allows us to achieve 3×3$\times$3 × speedup over the non-approximated LM optimizer, while maintaining similar performance, as shown in Table 3.
4.5 Learning Rate Scheduler
When second-order optimizers are applied to large-scale systems, numerical errors can occur, causing the solution vector $\Delta\boldsymbol{\beta}$ to overshoot the true loss landscape. To determine the optimal learning rate, line search algorithms or trust region methods [2, 18, 37, 11, 41, 7] are usually employed. However, these methods require additional forward or backward passes, increasing the computational overhead and slowing down the overall algorithm. Rather than relying on these methods, we estimate the learning rate by constraining the maximum update of color values. In 3DGS, the color parameters range between $\approx-1.77$ and $\approx 1.77$ due to the level 00 spherical harmonic coefficient. This gives us a natural bound for the color parameters. We trust the update direction $\Delta\boldsymbol{\beta}$, as long as the resulting color change does not exceed 1. If the change surpasses this threshold, we scale $\Delta\boldsymbol{\beta}$ so that the maximum change remains 1. Notably, this scaling is applied uniformly across all parameters, regardless of their type (e.g., opacity, color, mean, scale, or rotation). The effectiveness of this learning rate heuristic is demonstrated in Table 4. Additionally, the supplementary document provides examples of learning rates assigned at each iteration.
Refer to caption Refer to caption Refer to caption
Figure 4: We plot the test set performance of baseline methods evaluated on the Lego scene. Our method achieves comparable metrics in significantly less time. All optimizers use the MSE loss function.
5 Results
Datasets and Metrics. We conduct our main experiments on the synthetic NeRF datasets [35], which include $8$ different scenes at $800\times 800$ resolution, each having 100 training and 200 testing camera views. To initialize the Gaussians, we follow 3DGS [21] and position $10{,}000$ Gaussians at random locations with random colors inside of the cube encapsulating the object. We also report results on Tanks &\&& Temples [24], which includes $2$ scenes: a truck, and a train scene. In these datasets, the Gaussians are initialized with structure-from-motion point clouds, and the number of Gaussians is $136{,}029$ and $182{,}686$, respectively. For all experiments, we use the default training and test splits. We report the test set performance based on structural similarity index (SSIM) [51], learned perceptual image patch similarity score (LPIPS) [56], and peak signal-to-noise ratio (PSNR).
Baselines. We compare our optimizer against Adam [23], RMSprop [50], and SGD with momentum [45], using their respective implementations in PyTorch. The momentum term of SGD and second-moment factor of RMSprop is set to $0.99$. We used the default learning rates given in 3DGS [21] for Adam and RMSProp optimizers. For SGD, we used the learning rates $0.16$ for the location, $0.2$ for color and $0.1$ for other parameters. We also apply a learning rate schedule for the mean value parameters, as suggested in 3DGS. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [30], which enhances the efficiency of the backward computation by storing the gradient state at every $32^{\text{nd}}$ splat. All baselines are run for $10{,}000$ iterations.
Implementation Details. All experiments are performed on a single NVIDIA Tesla A40 GPU. We use a uniform damping parameter $\lambda=0.01$ across all experiments. Our method, as well as all baselines, optimize using the mean squared error (MSE) loss function. We run our method for $200$ iterations, setting the batch size to $8$ and the number of sampled pixels per tile ($N$) to $32$, using our loss-based sampling distribution $q_{\text{loss}}$. Maximum conjugate gradient iterations start from $3$ and increase to $8$ after the $50^{th}$ iteration.
5.1 Comparison
Next, we evaluate our method, named as Levenberg-Marquardt with Importance Sampling (LM-IS), on synthetic NeRF datasets and present the results in Table 1. Our approach achieves strong performance while significantly reducing computation time. We report average metrics across all the synthetic NeRF datasets. Additionally, Figure 4 illustrates the test set performance on the Lego scene over time. As shown, our second-order optimizer converges rapidly and achieves competitive results across all metrics. Overall, our method demonstrates substantial improvements over the baselines in both efficiency and accuracy.
Another observation is that SGD with momentum consistently underperforms compared to other optimizers, despite our efforts at hyperparameter tuning. In fact, we had to truncate its training process to fit its results within the same plot as other optimizers. This highlights the importance of adaptive learning rates for optimizing 3DGS, a feature incorporated in all other evaluated optimizers. We also did not observe significant speed benefits from the optimizations of Taming 3DGS, which is likely due to a comparatively lower number of Gaussians in this setting that reduces the advantages of gradient caching.
Please refer to the supplemental document for the results using SSIM loss, as well as results on the Tanks &\&& Temples dataset.
Table 1: We report LPIPS, SSIM and PSNR scores averaged across NeRF synthetic test datasets. Our method achieves similar quality in a shorter amount of time.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
3DGS-SGD 0.331 0.488 18.56 164.2
3DGS-RMSprop 0.159 0.856 25.52 54.42
3DGS-Adam 0.162 0.855 25.49 58.46
Taming3DGS-Adam 0.147 0.862 26.01 60.73
LM-IS (Ours) 0.144 0.867 25.82 12.76
5.2 Ablation Studies
In this section, we ablate various design choices incorporated into our optimizer. All of the results in this section are obtained by averaging the metrics across all datasets.
View Sampling. We propose a view sampling approach in Sec. 4.3 to ensure that our optimizer can effectively estimate the local curvature of the loss landscape while still maintaining a relatively low batch size, through sampling of diverse views in a batch. In Tab. 2, we ablate our design choice, where we observe that our method shows improvements as compared to random view sampling. Furthermore, we also report results on lower batch sizes, where the performance is significantly lower, further verifying that second-order optimizers indeed require larger batch sizes than first-order optimizers.
Table 2: Our optimizer relies on view sampling to obtain meaningful Jacobian approximations. If the number of images in the batch is limited, or if we use random sampling of images, the performance of the optimizer drops.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
Cluster - Batch Size = 1 0.372 0.153 13.11
Cluster - Batch Size = 2 0.329 0.472 15.65
Cluster - Batch Size = 4 0.174 0.841 23.97
Random - Batch Size = 8 0.150 0.863 25.45
Cluster - Batch Size = 8 0.144 0.867 25.82
Importance Sampling. Our proposed importance sampling approach in Sec. 4.4 is the core algorithm that provides significant speed benefits, enabling faster processing by effectively reducing the dimensionality of highly redundant Jacobians. Here, we compare our performance against three baselines in Tab. 3. Firstly, when no sampling is used at all, i.e., the full Jacobians are considered every time, the amount of training time required increases significantly, showing the importance of the sampling and approximation. When we adopt our importance sampling but a high number of sampled pixels per tile (e.g., $N=128$) is used, there are some gains over no sampling, but its efficiency is still suboptimal. Furthermore, when we adopt a simple random sampling ($q_{\text{uniform}}$) instead of our importance sampling, we also find that efficiency and accuracy drops. All these show the efficacy of our importance sampling method.
Table 3: This table illustrates the effectiveness of our loss-based sampling distribution. Even when the number of samples per tile ($N$) is low, we can effectively estimate the loss function and reach a similar performance that we would get without sampling. We also show that importance sampling performs slightly better than uniform sampling on average.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
LM - No Sampling 0.143 0.868 25.88 37.70
LM - $q_{\text{loss}}$ $N=128$ 0.142 0.868 25.91 24.20
LM - $q_{\text{uniform}}$ $N=32$ 0.147 0.866 25.68 13.16
LM - $q_{\text{loss}}$ $N=32$ 0.144 0.867 25.82 12.76
Learning Rate Heuristic. We compare our learning rate heuristic introduced in Sec. 4.5 against uniform learning rate, Armijo line search [2, 41], and a grid search method in Tab. 4.
The grid search method selects the learning rate at each iteration by identifying the value that results in the greatest reduction in the objective function. Although this method is stable, the learning rates obtained are too pessimistic, and it cannot reach the quality of our heuristic. Armijo line search, or backtracking line search, improves the grid search method by picking the highest learning rate that results in sufficient reduction of the loss. While this approach has been successfully integrated as an improvement in deterministic second-order optimizers, our findings indicate that its effectiveness is limited in stochastic settings, such as ours, where optimization relies on multiple approximations. The performance of the line search methods can be tuned by using fewer candidate learning rates, however, we found that even with a large set of candidates, they fail to reach the quality of our heuristic within the same number of iterations.
Table 4: We compare the effect of various learning rate schedules. Our learning rate heuristic is able to assign dynamic learning rates resulting in rapid convergence rate and stability.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
Uniform $LR=0.5$ 0.504 0.365 9.46 18.32
Uniform $LR=0.3$ 0.380 0.544 13.64 15.57
Uniform $LR=0.1$ 0.161 0.855 25.37 14.33
Grid Line Search 0.262 0.756 21.64 761.7
Armijo Line Search 0.363 0.658 13.34 51.75
Our Heuristic (Sec. 4.5) 0.144 0.867 25.82 12.76
6 Discussion and Conclusion
Limitations. While our method is computationally efficient for a lower number of Gaussians, its advantages diminish as the number increases, which may limit scalability in highly complex scenes. Exploring importance sampling within the local neighborhood of the Gaussians could be an interesting research direction to address this limitation. Moreover, we employ mean squared error instead of mean absolute error and only used a diagonal approximation of the original SSIM loss, unlike the original Adam-based optimization for 3D Gaussian Splatting. This discrepancy in loss formulations may affect convergence behavior, but future work could explore re-weighting strategies of residuals to obtain different norms.
By leveraging the inherent sparsity of the Jacobian matrix and integrating a GPU-parallelized conjugate gradient solver, our method significantly reduces both memory consumption and computational overhead. Our novel view and pixel-wise importance sampling further enhance efficiency, enabling rapid convergence by decreasing the per-step overhead. Additionally, our heuristic for learning rate selection eliminates the need for costly line search procedures, further accelerating training. Our approach achieves up to a 6× speedup over Adam, particularly excelling in scenarios with a low number of Gaussians. Overall, our results highlight the potential of second-order methods in accelerating optimization for 3D Gaussian Splatting. We anticipate that future work will refine these techniques, particularly in handling larger Gaussian counts and incorporating additional perceptual loss terms, further advancing the efficiency and quality of Gaussian-based scene representations.
Supplementary Material
Appendix A Details of CUDA Implementation
We follow the parallelization pattern applied in 3DGS [21], dividing images into $16\times 16$ blocks, and launching threads for each pixel. Our main difference from 3DGS is that our block size is significantly less than $16\times 16=256$, due to pixel sampling. We set the number of pixels sampled per block as multiples of $32$, which allows us to do fast warp level reductions when needed. Now, we describe the Jacobian-vector product kernels in detail:
Jv Kernel. This kernel is similar to backward pass of 3DGS, and can be split into two kernel calls. We launch a thread for sampled columns (which corresponds to pixels) of $\mathbf{J}^{\top}$, and loop over the Gaussians. However, when the sample size is small, i.e 32, only 32 threads (or one warp per block) are launched. The limited number of threads does not fully utilize the GPU’s resources, leading to inefficiencies. To mitigate, we launch additional threads along the Gaussian direction. Each thread processes part of the Gaussian loop, skipping most calculations except for four dependent variables (transmittance and accumulated RGB values). These values are simply recalculated by each thread. Warp level reductions then finalize the dot product.
Jv Kernel. This kernel also uses per pixel parallelization and computes the dot product. However, it cannot be split into two smaller parts as the $\mathbf{J}^{\top}\mathbf{v}$ kernel. Therefore, we utilize the shared memory more by bringing additional Gaussian parameters. Warp-level reductions are not needed; each thread independently computes and writes its local sum.
DiagJJ Kernel. Similarly, this kernel cannot be split, so we again leverage shared memory to hold more data. As we compute derivatives, we square and accumulate them using warp-level reductions.
For the rest of the conjugate gradient algorithm, we follow the implementation described in [52]. This version requires fewer kernel calls when compared to a naive implementation and has been successfully applied in other works [16, 20, 57]. For the details, please see the respective papers.
Table 5: Comparisons on Tanks &\&& Temples scenes. All optimizers use MSE loss.
[…]
3DGS-Adam Train 0.321 0.670 20.12 130.0
LM-IS (Ours) 0.302 0.694 20.16 145.28
3DGS-Adam Truck 0.185 0.780 23.44 128.18
LM-IS (Ours) 0.203 0.772 22.93 119.83
3DGS-Adam Avg. 0.253 0.725 21.78 129.10
LM-IS (Ours) 0.253 0.733 21.55 132.56
Appendix B Results Per Scene
We share the per-scene scores obtained in NeRF synthetic dataset [35] in Table 6. In this experiment, all of the optimizers use MSE loss.
Table 6: Comparisons of different methods on the synthetic NeRF dataset. All optimizers use MSE loss.
Method Scene LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\downarrow$ Time (s)
3DGS-RMSprop Chair 0.162 0.865 24.96 54.11
3DGS-Adam 0.187 0.834 22.09 60.22
Taming3DGS-Adam 0.145 0.876 25.75 61.32
LM-IS (Ours) 0.147 0.874 25.39 13.53
3DGS-RMSprop Drums 0.185 0.834 22.05 52.73
3DGS-Adam 0.164 0.864 24.92 55.36
Taming3DGS-Adam 0.166 0.845 22.57 52.82
LM-IS (Ours) 0.166 0.845 22.30 12.26
3DGS-RMSprop Ficus 0.090 0.893 25.79 56.21
3DGS-Adam 0.093 0.890 25.60 56.92
Taming3DGS-Adam 0.078 0.903 26.51 60.74
LM-IS (Ours) 0.094 0.891 25.55 10.01
3DGS-RMSprop Hotdog 0.111 0.923 29.65 53.72
3DGS-Adam 0.111 0.924 29.78 57.85
Taming3DGS-Adam 0.099 0.929 30.34 60.99
LM-IS (Ours) 0.088 0.933 30.55 13.29
3DGS-RMSprop Lego 0.196 0.814 24.34 53.55
3DGS-Adam 0.201 0.809 24.23 57.73
Taming3DGS-Adam 0.184 0.821 24.82 59.97
LM-IS (Ours) 0.180 0.827 24.88 13.88
3DGS-RMSprop Materials 0.182 0.826 23.95 51.48
3DGS-Adam 0.183 0.827 23.95 56.11
Taming3DGS-Adam 0.167 0.829 24.22 62.11
LM-IS (Ours) 0.158 0.862 24.21 11.73
3DGS-RMSprop Mic 0.104 0.922 28.14 50.48
3DGS-Adam 0.108 0.920 28.09 56.02
Taming3DGS-Adam 0.100 0.925 28.64 63.34
LM-IS (Ours) 0.093 0.930 28.10 11.80
3DGS-RMSprop Ship 0.238 0.770 25.28 63.03
3DGS-Adam 0.274 0.769 25.27 67.41
Taming3DGS-Adam 0.238 0.767 25.44 64.49
LM-IS (Ours) 0.225 0.784 25.59 15.57
3DGS-RMSprop Average 0.159 0.856 25.52 54.42
3DGS-Adam 0.162 0.55 25.49 58.46
Taming3DGS-Adam 0.147 0.862 26.01 60.73
LM-IS (Ours) 0.144 0.867 25.82 12.76
We also provide experiments with SSIM loss, which is approximated diagonally, as mentioned in the main paper. The original SSIM loss propagates gradients through the local neighbors, which is not efficient with per-pixel parallelization scheme of our $\mathbf{Jv}$ kernel. In Table 7, we show that our optimizer still outperforms the competing methods even with the diagonal approximation. All optimizers use a weighted average of MSE and SSIM losses, with weights $1.0$ and $0.2$, respectively.
Table 7: Comparisons of different methods on the synthetic NeRF dataset. All optimizers use a weighted average of MSE and SSIM loss, with weights $1.0$ and $0.2$ respectively.
Method Scene LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\downarrow$ Time (s)
3DGS-RMSprop Chair 0.155 0.878 24.88 85.70
3DGS-Adam 0.157 0.878 24.88 87.42
Taming3DGS-Adam 0.138 0.890 25.66 86.58
LM-IS (LM-IS (Ours)) 0.135 0.888 25.88 38.93
3DGS-RMSprop Drums 0.190 0.853 21.52 83.96
3DGS-Adam 0.190 0.853 21.60 86.46
Taming3DGS-Adam 0.159 0.872 22.38 86.82
LM-IS (LM-IS (Ours)) 0.152 0.870 22.69 36.68
3DGS-RMSprop Ficus 0.089 0.901 25.34 84.59
3DGS-Adam 0.090 0.901 25.41 83.50
Taming3DGS-Adam 0.074 0.913 26.28 87.88
LM-IS (LM-IS (Ours)) 0.082 0.907 25.85 33.08
3DGS-RMSprop Hotdog 0.098 0.937 29.47 86.06
3DGS-Adam 0.096 0.938 29.64 85.20
Taming3DGS-Adam 0.087 0.942 30.29 84.79
LM-IS (LM-IS (Ours)) 0.078 0.945 31.28 36.92
3DGS-RMSprop Lego 0.189 0.835 24.15 83.55
3DGS-Adam 0.192 0.836 24.170 86.41
Taming3DGS-Adam 0.175 0.847 24.68 85.47
LM-IS (Ours) 0.165 0.846 25.27 43.77
3DGS-RMSprop Materials 0.175 0.864 23.84 81.72
3DGS-Adam 0.176 0.863 23.80 85.35
Taming3DGS-Adam 0.159 0.870 24.12 85.55
LM-IS (Ours) 0.146 0.871 24.28 41.16
3DGS-RMSprop Mic 0.103 0.936 27.90 81.66
3DGS-Adam 0.109 0.933 27.75 80.23
Taming3DGS-Adam 0.090 0.941 28.30 84.91
LM-IS (Ours) 0.083 0.942 28.48 34.26
3DGS-RMSprop Ship 0.227 0.804 22.77 86.02
3DGS-Adam 0.225 0.806 24.90 88.33
Taming3DGS-Adam 0.214 0.811 25.15 85.38
LM-IS (Ours) 0.219 0.806 25.71 46.40
3DGS-RMSprop Average 0.153 0.876 25.23 84.16
3DGS-Adam 0.154 0.876 25.27 85.37
Taming3DGS-Adam 0.137 0.886 25.86 85.93
LM-IS (Ours) 0.132 0.885 26.18 38.90
Appendix C Results in Real-world Datasets
We share our experiment results on the Tanks &\&& Temples dataset [24]. In this dataset, the initial number of Gaussians is larger than $100{,}000$. Although the performance of our optimizer drops, we can still compete with the Adam optimizer as shown in Table 5 and Fig 6. In this dataset, Adam optimizer is run for $10{,}000$ iterations, while ours is run for $130$ iterations. We use MSE loss for both of the optimizers.
Appendix D Example of Learning Rate Schedule
Figure 5: Our learning rate heuristic can assign both high and low learning rates dynamically, resulting in stable and rapid convergence.
Our learning rate scheduler can assign both low and high learning rates, while providing stable training. We share an example of learning rates obtained in Lego scene in Fig. 5. Note that we cap the maximum learning rate to $0.5$ in our experiments.
Truck Scene Train Scene
3DGS-Adam
Refer to caption Refer to caption Refer to caption Refer to caption
LM-IS (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: We show qualitative results on the real-world dataset. Our second-order optimizer reaches a similar quality in a similar amount of time.